OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin(7pi/8)

OpenStudy (anonymous):

Would I be using the formula: \[\sin \frac{ x }{ 2 }=\pm \sqrt{\frac{ 1+cosx }{ 2 }}\]

OpenStudy (anonymous):

What exactly does that mean?

OpenStudy (jdoe0001):

hmmm what ?

OpenStudy (anonymous):

Those are half-angle formulas for sine and cosine. I don't understand what you're trying to imply.

OpenStudy (jdoe0001):

ohhh that's what you'd be using, is a -cos, not a +cos like you post it anyhow... \(\bf \cfrac{7}{8}\cdot 2\implies \cfrac{7}{4}\qquad \cfrac{7\pi}{8}\cdot 2\implies \cfrac{7\pi}{4}\\ \quad \\ sin\left(\frac{7\pi}{8}\right)\implies sin\left(\cfrac{\frac{7\pi}{4}}{2}\right)=\pm \sqrt{\cfrac{1-cos\left(\frac{7\pi}{4}\right)}{2}}\)

OpenStudy (jdoe0001):

and you can find \(\bf \cfrac{7\pi}{4}\) in your Unit Circle

OpenStudy (anonymous):

Oh, yeah sorry it was a typo! Thank you!

OpenStudy (jdoe0001):

yw